# TRACE Back from the Future: A Probabilistic Reasoning Approach to Controllable Language Generation

### v1

@@ Appendix A Factorizability of Attributes
The results indicate that both Toxicity and Politics exhibit non-factorizable characteristics to varying degrees, as expected for complex semantic attributes. Nonetheless, as discussed in Section 3, TRACE achieves strong empirical performance on these tasks even with the simpler factorized classifier, highlighting the effectiveness of combining it with exact EAP computation over the HMM.
Appendix B HMM Implementation Details
We employ standard Hidden Markov Models (HMMs) for all TRACE experiments, distilled from the base language models (GPT2-large, Gemma-2B) following the approach described in Zhang et al. (2023); Liu et al. (2024b).
@@ Appendix B HMM Implementation Details
Text sequences were sampled unconditionally from the respective base LMs. The HMM hidden state size was set to $h=4096$. Training sample sizes were:
@@
  • Gemma-2B HMM (Low-Data Ablation, HMM): 1 million samples.
@@
Appendix C Attribute Classifier Implementation Details
@@
As RealToxicityPrompts rarely elicit political content, training data was sourced from the News Category dataset (Misra, 2022). Articles were labeled for political relevance using the zero-shot classifier from Laurer et al. (2023), providing target probabilities $p$. We modeled non-politicalness (1-p) and fitted the log-linear weights using log-MSE loss against these scores after applying the logit transformation (Section 4.3) with scale $b=1$ and shift $c=-10$.
Appendix D Baseline Details
@@
Training times (Table 5) and inference time ratios (Table 6) for baselines were also sourced:
@@
Appendix E Evaluation Details

### v2

@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Factorizability of Attributes
The results indicate that both Toxicity and Politics exhibit non-factorizable characteristics to varying degrees, as expected for complex semantic attributes. Nonetheless, as discussed in Section 5.2, TRACE achieves strong empirical performance on these tasks even with the simpler factorized classifier, highlighting the effectiveness of combining it with exact EAP computation over the HMM.
Appendix B Additional Detoxification Metrics
To further support the findings in Section 5.2, we provide two supplementary evaluations.
Conditional Entropy.
Table 7 reports the conditional entropy of continuations given a prompt under each model. Lower entropy indicates less lexical and structural diversity, often symptomatic of repetitive or degenerate outputs. TRACE achieves entropy comparable to the base LM and much higher than DPO, reinforcing that RL methods tend to reduce generation diversity.
Method Entropy (↑)
GPT2-large 52.06
DPO 39.52
TRACE 52.54
GPT-4 LM-as-a-Judge Evaluation.
Table 8 presents human-aligned evaluations from GPT-4, comparing continuations across nontoxicity, fluency, and diversity. TRACE matches or exceeds DPO on nontoxicity and diversity, while maintaining comparable fluency to both Gemma2B and DPO, affirming that TRACE achieves strong controllability without harming output quality.
Table 8: GPT4 LM-as-a-judge Evaluations.
Method Nontoxicity (↑) Fluency (↑) Diversity (↑)
Gemma2B 4.39 3.76 2.93
DPO 4.65 3.94 2.86
TRACE 4.69 3.72 2.94
Impact of HMM Quality
The quality of the distilled Hidden Markov Model (HMM) has a direct impact on the effectiveness of TRACE for controllable generation. To illustrate this, Figure 5 plots the HMM’s log-likelihood on a validation set against the average maximum toxicity of the generations it produces at different points during its training. The results clearly show that as the HMM becomes a better probabilistic model of the language (higher log-likelihood), its ability to guide the generation towards non-toxic outputs improves (lower toxicity). This suggests that further advancements in HMM distillation techniques will likely yield even better performance from TRACE.
Figure 5: Effect of HMM Quality on Detoxification. As the HMM is trained for more steps, its fit to the data improves (increasing Log-Likelihood, blue), which corresponds to a decrease in the average maximum toxicity of generated text (red). This shows that higher-quality HMMs lead to more effective control.
Appendix C HMM Implementation Details
We employ standard Hidden Markov Models (HMMs) for all TRACE experiments, distilled from the base language models (GPT2-large, Gemma-2B) following the approach described in Zhang et al. (2023); Liu et al. (2023).
@@ Appendix C HMM Implementation Details
Text sequences were sampled unconditionally from the respective base LMs. We used two HMM configurations depending on the task. For the primary detoxification and composition tasks, the HMM hidden state size was set to $h=4096$. For the low-resource personalization experiments (Section 5.3), a smaller HMM with $h=256$ was used to demonstrate TRACE’s effectiveness even with a more compact model.
Training sample sizes for the larger ($h=4096$) HMM were:
@@
Training Details.
We initialize the parameters of the HMM using the latent variable distillation technique (Liu et al., 2023). For training of the HMM, we employed a mini-batch variant of expectation maximization (EM) (Sato, 1999; Peharz et al., 2020) which interpolates between the old and new parameters (based on a mini-batch) using a step size $\alpha$. We use a mini-batch size of 4096 and train for 50 epochs, and anneal the step-size according to a linear decay schedule from $1.0$ to $0.0$ following Zhang et al. (2025).
@@
Appendix D Attribute Classifier Implementation Details
@@
As RealToxicityPrompts rarely elicit political content, training data was sourced from the News Category dataset (Misra, 2022). Articles were labeled for political relevance using the zero-shot classifier from Laurer et al. (2023), providing target probabilities $p$. We modeled non-politicalness (1-p) and fitted the log-linear weights using log-MSE loss against these scores after applying the logit transformation (Section 4.3) with scale $b=1$ and shift c=10c=-10.
Appendix E Baseline Details
@@
Training times and inference time ratios (Table 4) for baselines were sourced from:
@@
Appendix F Evaluation Details